OpenStudy (anonymous):

factor each polynomial x^3+8x^2+15x

OpenStudy (anonymous):

do u no how to do this i need to know the steps

OpenStudy (anonymous):

0,-5,-3

OpenStudy (solomonzelman):

Factor out of ` x ` and then, Hint: ` 5 + 3 = 8 `

OpenStudy (anonymous):

x(x^2+5x+3x+15)=0 x=0,(x(x+5)+3(x+5))=0 x=0.x=-5,x=-3

OpenStudy (anonymous):

thank yall so much

OpenStudy (anonymous):

this the next one in my homeweork x^3+3x^2-2x-6